OpenStudy (anonymous):

Solve for m. -4=m/4-7

OpenStudy (anonymous):

@Abhisar

OpenStudy (nincompoop):

TRY TO GET RID OF THE DENOMINATOR FIRST ON THE RIGHT SIDE

OpenStudy (anonymous):

so get rid of -7?

OpenStudy (anonymous):

7*

OpenStudy (anonymous):

oh nvm 4

OpenStudy (nincompoop):

I will show you an example \(a = \large \frac{b}{c} \) to solve for b, we multiply both sides by c, \( a \times c = \large \frac{b}{c} \times c \) notice how \( \large \frac{1}{c} \times c = 1 \)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it would be -4 on both sides?

OpenStudy (nincompoop):

then we get \(ac = 1 \times b \rightarrow ac = b \)

OpenStudy (anonymous):

okay

OpenStudy (nincompoop):

well you need to simplify the denominator first!

OpenStudy (anonymous):

4

OpenStudy (nincompoop):

what is 4 - 7 ?

OpenStudy (anonymous):

um -3

OpenStudy (anonymous):

@nincompoop?

OpenStudy (nincompoop):

okay then solve for m

OpenStudy (anonymous):

so -7?

OpenStudy (nincompoop):

how?

OpenStudy (anonymous):

-3-4....

OpenStudy (anonymous):

i have no idea what i am doing

OpenStudy (anonymous):

oh wait it is 12!

OpenStudy (anonymous):

because you have to add 7 to both sides, then you simplify and then multiply like this: 3*4 which gives you 12...right?

OpenStudy (anonymous):

so 12=m?

OpenStudy (anonymous):

thx for the help if that was right ;)